# Rethinking Spectral Graph Neural Networks with Spatially Adaptive Filtering

### v3

@@ 3. Notations and Preliminaries
Figure 1. Distributions of connected nodes in the new graph based on their geodesic/shortest-path distance (as $\Delta_{i,j}$) in the original graph. Nodes, distant in the original graph ($\Delta_{i,j}$ 1 in x-axis), can be linked in the new graph (Number ¿ 0 in y-axis).
Figure 2. Left y-axis: Homophily comparison between original and new graphs, considering only positive edges (blue and yellow bars). Right y-axis: Percentage of edges connecting nodes from different classes, identified by negative edges (green bar).
4. Rethinking Spectral GNNs from the Spatial Perspective
@@
4.1. Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
The graph signal denoising problem (Shuman et al., 2013) was initially leveraged in (Ma et al., 2021; Zhu et al., 2021) as a means to interpret GNNs with smoothness assumption, which yet does not always hold in certain real-world graph scenarios such as heterophily (Zhu et al., 2020b). Without loss of generality, in this work, we consider a more generalized graph optimization problem [^1] This problem was first introduced in (He et al., 2021) for theoretically grounded graph filters. However, in this study, we repurpose it as a bridge between spectral filtering and spatial aggregation.
(3) $\mathop{\arg\min}_{\mathbf{Z}}\ \mathcal{L}=\alpha\|\mathbf{X}-\mathbf{Z}\|^{2}_{F}+(1-\alpha)\cdot\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$
where $\mathbf{Z}\in\mathbb{R}^{N\times d}$ refers to node representations, $\gamma_{\theta}(\hat{\mathbf{L}})$ determines the rate of propagation (Spielman, 2012) by operating on the graph spectrum, i.e., $\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{U}\gamma_{\theta}({\mathbf{\Lambda}})\mathbf{U}^{T}$, and $\alpha\in(0,1)$ is a trade-off coefficient. In case of setting $\gamma_{\theta}(\hat{\mathbf{L}})=\hat{\mathbf{L}}$, Eq. (3) turns into the well-known graph signal denoising problem. To ensure the convexity of the objective in Eq. (3), a positive semi-definite constraint is imposed on $\gamma_{\theta}(\hat{\mathbf{L}})$, i.e., $\gamma_{\theta}(\lambda)\geq 0$ for $\lambda\in\left[0,2\right]$. Then, one can address this minimization problem through either closed-form or iterative solutions.
Closed-form Solution. The closed-form solution can be obtained by setting the derivative of the objective function $\mathcal{L}$ to 0, i.e., $\frac{\partial\mathcal{L}}{\partial\mathbf{Z}}=2\alpha(\mathbf{Z}-\mathbf{X})+2(1-\alpha)\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z}=0$. Let $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$, we can observe that the closed-form solution in Eq. (4) is equivalent to the spectral filtering in Eq. (1).
(4) $\mathbf{Z}^{*}=(\mathbf{I}+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\hat{\mathbf{L}}))^{-1}\mathbf{X}=g_{\psi}(\hat{\mathbf{L}})\mathbf{X}=\mathbf{U}g_{\psi}(\mathbf{\Lambda})\mathbf{U}^{T}\mathbf{X}.$
@@ 4.1. Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
Iterative Solution. Alternatively, we can take an iterative gradient descent method such that $\mathbf{Z}^{(k)}=\mathbf{Z}^{(k-1)}-b\frac{\partial\mathcal{L}}{\partial\mathbf{Z}}|_{\mathbf{Z}=\mathbf{Z}^{(k-1)}}$ with a step size $b=\frac{1}{2}$, which yields a concise iterative solution in Eq. (5) with $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})$. Notably, by taking $\hat{\mathbf{A}}^{\text{new}}$ as a new computation graph, this solution closely mirrors the spatial aggregation in Eq. (2).
(5) $\mathbf{Z}^{(k)}=\alpha\mathbf{X}+(1-\alpha)\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K$
Theoretical Interaction — the Adapted New Graph. With the non-negative constraint, it is evident that both spectral filtering and spatial aggregation effectively address the generalized graph optimization problem in Eq. (3), despite their distinctive forms and operation domains. Upon closer examination, we discover a compelling relationship between the graph filter $g_{\psi}(\lambda)$ in Eq. (4) and the new graph $\hat{\mathbf{A}}^{\text{new}}$ in Eq. (5), given $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$,
(6) $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})$
@@
What are the differences between $\hat{\mathbf{A}}^{\text{new}}$ and $g_{\psi}(\hat{\mathbf{L}})$? Whereas the former as the uncovered new graph elucidates the inherent spatial node relationships, the latter is a graph operation, primarily processing graph features within spectral domain. It is crucial to understand that $g_{\psi}(\hat{\mathbf{L}})$ may not result in a dense matrix, especially with fixed-order polynomial approximation. This is because it captures up to only a $K$-hop neighborhood, i.e., $g_{\psi}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\psi_{k}P_{k}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\omega_{k}\hat{\mathbf{A}}^{k}$, practically limiting spectral GNNs’ effective propagation range. In contrast, our newfound graph $\hat{\mathbf{A}}^{\text{new}}$ intrinsically enjoys a non-local property, as confirmed in the following section. Building upon this discovery, we further devise a framework to break domain barriers, overcoming the limitations of current spectral GNNs due to truncated polynomials (see details in Section 5).
4.2. In-depth Analysis of the adapted new graph
@@
4.2.1. Non-locality
@@
With the established non-negative constraint on graph filters, specifically $0<g_{\psi}(\lambda)\leq 1$, it becomes evident that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ falls into the interval permitting Neumann series expansion, as shown in lemma 1 (Horn and Johnson, 2012). Building on this observation, we present a non-trivial property of the new graph in the following proposition (see proof in Appendix A.1).
Proposition 0 ().
@@
This proposition implies that the new graph engenders immediate links between nodes that originally necessitate multiple hops for connection. To further underpin this theoretical claim, we analyze the general connection status on the new graph by BernNet (He et al., 2021), a spectral GNN adhering to the non-negative constraint. From Figure 1, it is apparent that nodes originally separated by multiple hops achieve direct connections in the new graph.
4.2.2. Signed Edge Weights — Discerning Label Consistency
Upon further scrutinizing the adapted new graph, we make a notable discovery that it readily accommodates both positive and negative edge weights. A more granular analysis in Figure 2 reveals that a considerable portion of positive edge weights are assigned to the same-class node pairs, enhancing graph homophily (exemplified by edge homophily ratio (Zhu et al., 2020b)). Conversely, edges parameterized with negative weights tend to bridge nodes with different labels. These findings demonstrate the newfound graph’s adeptness in discerning label consistency among nodes. To theoretically explain this phenomenon, we further present the proposition below (see proof in Appendix A.2):
Proposition 0 ().
@@ 4.2.2. Signed Edge Weights — Discerning Label Consistency
To summarize, our investigation into spectral GNNs in the spatial domain reveals that graph spectral filtering fundamentally alters the original graph, imbuing it with non-locality and signed edge weights that capture label consistency among nodes. These findings highlight the interpretable role of spectral GNNs in the spatial domain, and prompt us to rethink current spectral GNNs beyond the truncated polynomial filters.
5. Spatially Adaptive Filtering Framework
@@
Building on our discoveries, we re-evaluate the state-of-the-art spectral GNNs and put forth a paradigm-shifting framework, Spatially Adaptive Filtering (SAF), for joint exploitation of graph-structured data across both spectral and spatial domains (refer to Figure 3). SAF leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation, addressing enduring challenges in GNNs related to long-range dependencies and graph heterophily.
5.1. Non-negative Spectral Filtering
@@
The proposed SAF requires explicit computation of the newfound graph, as outlined in Eq. (6). This further necessitates the graph filter $g_{\psi}:\left[0,2\right]\to\mathbb{R}$ to satisfy the non-negative constraint from Eq. (3): $0\leq g_{\psi}(\lambda)\leq 1$. However, not all extant graph filters fulfill this prerequisite. For instance, the filter use by GCN (Kipf and Welling, 2017), $g_{\psi}(\lambda)=1-\lambda$, takes negative values when $\lambda>1$. In this research, we approximate the graph filter using Bernstein polynomials (Farouki, 2012), which are known for their non-negative traits (Powers and Reznick, 2000) and are essential in a preeminent spectral GNN, BernNet (He et al., 2021). For $g_{\psi}(\lambda)\leq 1$ part, we rescale Bernstein polynomials with the following proposition.
Proposition 0 ().
@@
Let $B_{k,K}(x)$ denote the Bernstein polynomial basis of index $k$ and degree $K$, which is defined as $B_{k,K}(x)=\binom{K}{k}(1-x)^{K-k}x^{k}$ for $x\in[0,1]$. Let $\psi_{k}$ denote the $k$-th coefficient of a polynomial $p(x)$ of degree $K$, where $p(x)=\sum_{k=0}^{K}\psi_{k}B_{k,K}(x)$ with $\psi_{k}\geq 0$ for all $k$. Then for all $x\in[0,1]$, we have $g_{\psi}(x)\leq\max\{\psi_{k}\}_{k=0}^{K}$.
Proposition 1 suggests that the Bernstein polynomial function attains its maximum value in $\psi_{\text{max}}=\max\{\psi_{k}\}_{k=0}^{K}$. Therefore, $g_{\psi}(\lambda)$ can be rescaled within $[0,1]$ by $\hat{g}_{\psi}(\lambda)=\frac{1}{\psi_{\text{max}}}\sum_{k=0}^{K}\psi_{k}B_{k,K}(\frac{\lambda}{2})$, enabling us to formulate the spectral filtering in SAF as
$\mathbf{Z}_{f}=\hat{g}_{\psi}(\hat{\mathbf{L}})f_{\varphi}(\mathbf{X})=\frac{1}{\psi_{\text{max}}}\sum_{k=0}^{K}\psi_{k}\frac{1}{2^{K}}\binom{K}{k}(2\mathbf{I}-\hat{\mathbf{L}})^{K-k}\hat{\mathbf{L}}^{k}f_{\varphi}(\mathbf{X})$
@@
where $f_{\varphi}(\cdot)$, a two-layer MLP, maps $\mathbf{X}$ from $F$ to $C$ dimensions using 64 hidden units, and $\{\psi_{k}\}_{k=1}^{K}$ are non-negative learnable parameters. Note that SAF also permits alternative implementations such as using Chebyshev polynomials (Hammond et al., 2011; Defferrard et al., 2016) for graph filter learning, enhancing models like ChebNetII (He et al., 2022) (see details in Appendix E.4).
5.2. Non-local Spatial Aggregation
Once acquiring a suitable spectral filter $\hat{g}_{\psi}(\lambda)$, we compute the adapted new graph as $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\tau(\mathbf{U_{m}}g_{\psi}(\mathbf{\Lambda_{m}})^{-1}\mathbf{U_{m}}^{T}-\mathbf{I})$ by Eq. (6) where $\tau=\frac{\alpha}{1-\alpha}$ is a scaling parameter and a partial eigendecomposition can be employed to obtain only $m$ extremal eigenvalues (Lanczos, 1950), producing a low-rank, robust structure for $\hat{\mathbf{A}}^{\text{new}}$. Equipped with this newfound graph, we proceed to perform non-local aggregation:
$\mathbf{Z}^{(l)}=(1-\eta)\mathbf{Z}^{(0)}+\eta\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(l-1)},\quad l=1,2,\cdots,L$

### v5

@@ III-B Spatial Aggregation
IV Rethinking Spectral Graph Neural Networks from the Spatial Perspective
@@
IV-A Interplay of Spectral and Spatial Domains through the Lens of Graph Optimization
The graph signal denoising problem [8] was initially leveraged in [23, 24] as a means to interpret GNNs with smoothness assumption, which yet does not always hold in certain real-world graph scenarios such as heterophily [46]. Without loss of generality, in this work, we consider a more generalized graph optimization problem [^1] This problem was first introduced in [13] for theoretically grounded graph filters. However, in this study, we repurpose it as a bridge between spectral filtering and spatial aggregation.
$\mathop{\arg\min}_{\mathbf{Z}}\ \mathcal{L}=\alpha\|\mathbf{X}-\mathbf{Z}\|^{2}_{F}+(1-\alpha)\cdot\mathrm{tr}(\mathbf{Z}^{T}\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z})$ (3)
where $\mathbf{Z}\in\mathbb{R}^{N\times d}$ refers to node representations, $\gamma_{\theta}(\hat{\mathbf{L}})$ determines the rate of propagation [58] by operating on the graph spectrum, i.e., $\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{U}\gamma_{\theta}({\mathbf{\Lambda}})\mathbf{U}^{T}$, and $\alpha\in(0,1)$ is a trade-off coefficient. In case of setting $\gamma_{\theta}(\hat{\mathbf{L}})=\hat{\mathbf{L}}$, Eq. (3) turns into the well-known graph signal denoising problem. To ensure the convexity of the objective in Eq. (3), a positive semi-definite constraint is imposed on $\gamma_{\theta}(\hat{\mathbf{L}})$, i.e., $\gamma_{\theta}(\lambda)\geq 0$ for $\lambda\in\left[0,2\right]$. Then, one can address this minimization problem through either closed-form or iterative solutions.
IV-A1 Closed-form Solution
The closed-form solution can be obtained by setting the derivative of the objective function $\mathcal{L}$ to 0, i.e., $\frac{\partial\mathcal{L}}{\partial\mathbf{Z}}=2\alpha(\mathbf{Z}-\mathbf{X})+2(1-\alpha)\gamma_{\theta}(\hat{\mathbf{L}})\mathbf{Z}=0$. Let $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$, we can observe that the closed-form solution in Eq. (4) is equivalent to the spectral filtering in Eq. (1).
$\mathbf{Z}^{*}=(\mathbf{I}+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\hat{\mathbf{L}}))^{-1}\mathbf{X}=g_{\psi}(\hat{\mathbf{L}})\mathbf{X}$ (4)
@@ IV-A1 Closed-form Solution
IV-A2 Iterative Solution
Alternatively, we can take an iterative gradient descent method such that $\mathbf{Z}^{(k)}=\mathbf{Z}^{(k-1)}-b\frac{\partial\mathcal{L}}{\partial\mathbf{Z}}|_{\mathbf{Z}=\mathbf{Z}^{(k-1)}}$ with a step size $b=\frac{1}{2}$, which yields a concise iterative solution in Eq. (5) with $\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})$. Notably, by taking $\hat{\mathbf{A}}^{\text{new}}$ as a new computation graph, this solution closely mirrors the spatial aggregation in Eq. (2).
$\mathbf{Z}^{(k)}=\alpha\mathbf{X}+(1-\alpha)\hat{\mathbf{A}}^{\text{new}}\mathbf{Z}^{(k-1)},\quad k=1,2,\cdots,K$ (5)
IV-A3 Theoretical Interaction — the Adapted New Graph
With the non-negative constraint, it is evident that both spectral filtering and spatial aggregation effectively address the generalized graph optimization problem in Eq. (3), despite their distinctive forms and operation domains. Upon closer examination, we discover a compelling relationship between the graph filter $g_{\psi}(\lambda)$ in Eq. (4) and the new graph $\hat{\mathbf{A}}^{\text{new}}$ in Eq. (5), given $g_{\psi}(\lambda)=(1+\frac{1-\alpha}{\alpha}\gamma_{\theta}(\mathbf{\lambda}))^{-1}$,
$\hat{\mathbf{A}}^{\text{new}}=\mathbf{I}-\gamma_{\theta}(\hat{\mathbf{L}})=\mathbf{I}-\frac{\alpha}{1-\alpha}(g_{\psi}(\hat{\mathbf{L}})^{-1}-\mathbf{I})$ (6)
@@ IV-A3 Theoretical Interaction — the Adapted New Graph
IV-A4 What are the differences between $\hat{\mathbf{A}}^{\text{new}}$ and $g_{\psi}(\hat{\mathbf{L}})$
Whereas the former as the uncovered new graph elucidates the inherent spatial node relationships, the latter is a operation in spectral GNNs that processes graph features within spectral domain. It is crucial to understand that $g_{\psi}(\hat{\mathbf{L}})$ may not result in a dense matrix, especially with fixed-order polynomial approximation. This is because it captures up to only a $K$-hop neighborhood, i.e., $g_{\psi}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\psi_{k}P_{k}(\hat{\mathbf{L}})=\sum_{k=0}^{K}\omega_{k}\hat{\mathbf{A}}^{k}$, practically limiting spectral GNNs’ effective propagation range. In contrast, our newfound graph $\hat{\mathbf{A}}^{\text{new}}$ intrinsically enjoys a non-local property, as confirmed in the following section. Building upon this discovery, we further devise a framework to transcend domain barriers, overcoming the limitations of current spectral GNNs due to truncated polynomials (see details in Section V).
IV-B In-depth Analysis of the Adapted New Graph
@@
Figure 1: Distributions of connected nodes in the new graph based on their geodesic/shortest-path distance (as $\Delta_{i,j}$) in the original graph. Nodes, distant in the original graph ($\Delta_{i,j}>1$ in x-axis), can be linked in the new graph (Number $>0$ in y-axis).
Figure 2: Left y-axis: Homophily comparison between original and new graphs, considering only positive edges (blue and yellow bars). Right y-axis: Percentage of negative edges in the new graph that connect nodes from different classes (green bar).
IV-B1 Non-locality
@@
With the established non-negative constraint on graph filters, specifically $0<g_{\psi}(\lambda)\leq 1$, it becomes evident that the eigenvalues of $\mathbf{I}-g_{\psi}(\hat{\mathbf{L}})$ falls into the interval permitting Neumann series expansion, as shown in lemma 1 [59]. Building on this observation, we present a non-trivial property of the new graph in the following proposition.
Proposition 1.
@@
Proof.
This proposition implies that the new graph engenders immediate links between nodes that originally necessitate multiple hops for connection. To further underpin this theoretical claim, we analyze the general connection status on the new graph by BernNet [13], a spectral GNN adhering to the non-negative constraint. From Fig. 1, it is apparent that nodes originally separated by multiple hops achieve direct connections in the new graph.
IV-B2 Signed Edge Weights — Discerning Label Consistency
Upon further scrutinizing the adapted new graph, we make a notable discovery that it readily accommodates both positive and negative edge weights. A more granular analysis in Fig. 2 reveals that a considerable portion of positive edge weights are assigned to the same-class node pairs, enhancing graph homophily (exemplified by edge homophily ratio [46]). Conversely, edges parameterized with negative weights tend to bridge nodes with different labels. These findings demonstrate the newfound graph’s adeptness in discerning label consistency among nodes. To theoretically explain this phenomenon, we further present the proposition below:
Proposition 2.
@@ IV-B3 Summarization
Our investigation into spectral GNNs in the spatial domain reveals that they fundamentally alters the original graph, imbuing it with non-locality and signed edge weights that capture label relationships among nodes. These findings highlight the interpretable role of spectral GNNs in the spatial domain and prompt us to rethink current models beyond the truncated polynomial filters.
Figure 3: Illustration of SAF framework where varying node colors represent different node labels. SAF leverages the adapted new graph by spectral filtering for auxiliary non-local aggregation in the spatial domain and allows individual nodes to flexibly balance between spectral and spatial features.
V Spatially Adaptive Filtering Framework
@@
Building on our discoveries, we re-evaluate the state-of-the-art spectral GNNs and put forth a paradigm-shifting framework, Spatially Adaptive Filtering (SAF), for joint exploitation of graph-structured data across both spectral and spatial domains. SAF leverages the adapted new graph by spectral filtering for an auxiliary non-local aggregation, addressing enduring challenges in GNNs related to long-range dependencies and graph heterophily. See the overall pipeline in Fig. 3.
V-1 Non-negative Spectral Filtering
@@
The proposed SAF requires explicit computation of the newfound graph, as outlined in Eq. (6). This further necessitates the graph filter $g_{\psi}:\left[0,2\right]\to\mathbb{R}$ to satisfy the non-negative constraint from Eq. (3): $0\leq g_{\psi}(\lambda)\leq 1$. However, not all extant graph filters fulfill this prerequisite. For instance, the filter use by GCN [10], $g_{\psi}(\lambda)=1-\lambda$, takes negative values when $\lambda>1$. In this research, we approximate the graph filter using Bernstein polynomials [60], which are known for their non-negative traits [61] and are essential in a preeminent spectral GNN, BernNet [13]. For $g_{\psi}(\lambda)\leq 1$ part, we rescale Bernstein polynomials with the following proposition.
Proposition 3.
@@
Let $B_{k,K}(x)$ denote the Bernstein polynomial basis of index $k$ and order $K$, which is defined as $B_{k,K}(x)=\binom{K}{k}(1-x)^{K-k}x^{k}$ for $x\in[0,1]$. Let $\psi_{k}$ denote the $k$-th coefficient of a polynomial $p(x)$ of order $K$, where $p(x)=\sum_{k=0}^{K}\psi_{k}B_{k,K}(x)$ with $\psi_{k}\geq 0$ for all $k$. Then for all $x\in[0,1]$, we have $g_{\psi}(x)\leq\max\{\psi_{k}\}_{k=0}^{K}$.
Proof.
$p(x)$ $=\sum_{k=0}^{K}\psi_{k}\binom{K}{k}(1-x)^{K-k}x^{k}$
$\leq\psi_{\text{max}}\sum_{k=0}^{K}\binom{K}{k}(1-x)^{K-k}x^{k}$
$=\psi_{\text{max}}(1-x+x)^{K}=\psi_{\text{max}}.$
Therefore, we have $p(x)\leq\max\{\psi_{k}\}_{k=0}^{K}$ for all $x\in[0,1]$. ∎
@@
Proposition 3 suggests that the Bernstein polynomial function attains its maximum value in $\psi_{\text{max}}=\max\{\psi_{k}\}_{k=0}^{K}$. Therefore, $g_{\psi}(\lambda)$ can be rescaled within $[0,1]$ by $\hat{g}_{\psi}(\lambda)=\frac{1}{\psi_{\text{max}}}\sum_{k=0}^{K}\psi_{k}B_{k,K}(\frac{\lambda}{2})$, which enables us to formulate the spectral filtering in SAF as $\mathbf{Z}_{f}=\hat{g}_{\psi}(\hat{\mathbf{L}})f_{\varphi}(\mathbf{X})=\frac{1}{\psi_{\text{max}}}\sum_{k=0}^{K}\psi_{k}\frac{1}{2^{K}}\binom{K}{k}(2\mathbf{I}-\hat{\mathbf{L}})^{K-k}\hat{\mathbf{L}}^{k}f_{\varphi}(\mathbf{X})$ where $f_{\varphi}(\cdot)$, a two-layer MLP, maps $\mathbf{X}$ from $F$ to $C$ dimensions using 64 hidden units, and $\{\psi_{k}\}_{k=1}^{K}$ are non-negative learnable parameters. Note that SAF also permits alternative implementations such as using Chebyshev polynomials [62, 9] for graph filter learning, enhancing models like ChebNetII [15] (see details in Section VI-H).
V-2 Non-local Spatial Aggregation